Sandra ran 7/12 of a mile. Lamar ran 3/4 of a mile. Who ran farther? Explain.
1
Expert's answer
2017-03-09T09:42:05-0500
3/4 = (3*3)/(4*3) = 9/12, but
9/12 > 7/12, hence 3/4 > 7/12.
Therefore, Lamar ran farther.